OpenStudy (anonymous):

14. The volume of a sphere is 5,000 pi m3. What is the surface area of the sphere to the nearest square meter?

OpenStudy (accessdenied):

Well, you know the volume. The surface area and volume are both equations that relate to the radius of the sphere. Maybe if you substitute \( V=5000 \pi \) and solve for r in that volume equation \( V = \frac{4}{3}\pi r^3 \), you can use it in the surface area equation \(A = 4\pi r^2 \).